[Brahms-l] Shift report 20040104 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 04 2004 - 00:08:44 EST
------------------------------------------------------------
Date: Sun Jan  4 00:08:44 2004, Shift: 16:00-24:00
Supervisor: sjs, with: 
Log book#15 pages 22 - 22

Shift Summary:
------------------------------------------------------------
1. Pathetic
2. pedestal run
3. none noticed

No useful beam until about 10 minutes to midnight. Although the
MCR claimed "physics running" much earlier, our background rates 
were very high until the get the collimators right, and that took
a while.
Pedestal runs 9345 and 9346
Started data run 9346 with FS at 1/4B, 8deg, MRS A1050 at 90 deg.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 4 00:09:06 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 04 2004 - 00:09:24 EST